Greenhab Report – March 30th

GreenHab Officer: Cheyenne Chamart

Environmental control:

Average temperatures: 85.2° F for the day (106° F around midday/in the afternoon)

Hours of supplemental light: N/A

Daily water usage for crops: N/A

Daily water usage for research and/or other purposes: N/A

Water in Blue Tank (full – not used) gallons

Time(s) of watering for crops: N/A

Changes to crops: N/A

Narrative: Today I stayed in the GreenHab for some time to water my plants (I used the tap water of the Hab as I only need 900 mL). I checked its, as well as the Science Dome, environmental parameters (t°, humidity, PAR,…) multiple times during the day. The GreenHab has a constant humidity of 9% during the day but is higher in the evening (18%).

Harvest: N/A

Support/supplies needed: Nothing

Journalist Report – March 30th

Crew 227 Journalist Report 30th March 2022
Sol: 3
Name of the person filing the report: Sirga Drouet
Sol 3 of the Tharsis mission!

Yesterday evening was a big tortilla feast thanks to Julien and Julie, the Juju, a small dedication to Ignacio for the perfect cooking, we enjoyed ourselves!

That evening we told each other anecdotes of our family trip. Small wink to our close relations: you made us laugh! Between Cheyenne who wanted to dive with her dad and Audrey who made a photo in China with a general of the army! Attention now we know your family secrets hehe! Then again the installation of polysomnographs (thank you CIDELEC), so this morning it was the sausages party in the community.

Everything passes so quickly we do not see the days passed, finally what reminds us of the notion of the time is the meals, especially the dinner which leads us to the moment to go to sleep! We went to bed at 1 am last night and we got up ready to start the day after 6 hours of sleep… Okay we abuse, we promise to try to come back from the mission without looking like aliens (what if we are on Mars? Are we aliens for you? Mmmh I let you choose 😉 ).

After a breakfast, which turned on the subject of the number of petals of cereals which we could put in our bowl per day, the surgical matches took again! Team mechanics against team anatomy for the moment the team mechanics stand out from all the teams! But we have more than one trick up our sleeve, we’ll end up knowing how to repair a fracture! In the end, our primary instinct gathered us around the lunch.

A loaf of good homemade bread with a touch of me to allow us to save flour. Bugs! To my surprise, everyone found the bread delicious! I don’t know if it’s because we are hungry all the time with all the activities or if we are just weird? Anyway, now that everybody likes this bread we’re going to eat it always with crushed flour worms ;-). I present you the brand new dish from the 5-star kitchen of Mars the Bugs’ Bread

And the adventure on the ground of Mars in all that you may ask?!

The 5 lucky ones to go on a mission are Audrey, Ignacio, Julie, Jean, and Julien.

Their mission, change Cyril’s antenna from Spirit and put Audrey’s bacteria in Martian exposure. The poor things are all scorching like crayfish hanging on a stone in front of the science dome. We can say hello to them from the window. For the antenna, the EVA team put the rover in front of the science dome while Cyril showed them exactly what to do. The thin team then left to explore the North Ridge area. RIP me Cyril and Cheyenne who have for super mission of the afternoon, the dishes…

However, being more than three in the station was good for us, it was calmer and after the last few days of being active at 200% without respite, slowing down is necessary for our bodies. This doesn’t mean that we did nothing! I accompanied Cheyenne for her sunshine and temperature measurements around her plants in different parts of the station. It is currently 42 degrees Celsius in the GreenHab (yes, we can use it to put Cheyenne’s seedlings). The time passed quickly and the adventurous team came back with a lot of discoveries to tell.

Small summary by Ignacio: After two days in the Hab this exit was saved, with the sun in the full, the soft sound of the voices on the headphones, according to him this expedition changed him deeply on his perception of the world and of himself. Finally, who is he? Olalala, what a poet! Nah I’m just kidding he’s not that poetic he just found this exit great, I wanted to add some romance but lost with him! Wait Ignacio may not be a poet but he is one hell of a 3D printer genius! Thanks to him we can use our gas stove without risking an explosion! A piece of quality and design, the dream of all gas stoves. Tonight we plan a Belgian dish, Belgian fries yeah!

At the moment we are in full medical visit and apparently, I do not breathe enough… Probably because tomorrow our revenge will sound and that I am super happy. I’m going to EVA with the kids who stayed in the station with me today!

On this, ciao!

See you tomorrow,
Sirga

Commander Report – March 30th

Crew 227 Commander Report 30th March 2022
SOL: 3
Name of person filing report: Cyril Wain

Sunny day today!

As usual, our day started at 8h00 with breakfast. This time we had breakfast with frosted flakes and dry milk. It was hard for the team to take only one cup of it. At the same time, Julien gave a look at the data of the polysomnographs taken during the night. Apparently, I snored 1453 times during the night. Hopefully, for my crew, I am isolated in my commander’s bedroom.

Around 09h00, we took some pictures of the team and other ones for some of our sponsors.

Between 11h00 and 12h00, Julie created a third crisis situation. This time, it was Julien and Audrey against Ignacio and me. Ignacio did really a great job there! On the other side, Julien was a bit disappointed with his result. One more point for the engineering team!!

At 12h, we had lunch. We ate some beans, peppers, tomatoes, carrots with some bread. This time the flour of the bread was replaced in part with flour of worms. The visual was identical to a classic bread and I have to say that the taste was quite similar as well. Everyone on the team appreciated it!

Between 13h00 and 16h00, Julien, Jean, Julie, Ignacio, and Audrey went out for an EVA. They started to work on my antenna in front of the science dome so I was able to guide them. They did an amazing job and I thank them so much for that! Then, they went around North Bridge to do some exploration and to place some samples for Audrey’s experiment.
During that time, Cheyenne, Sirga, and I worked on our personal experiment.

At 17h00, we gave a look at the pictures taken by the EVA crew of the day. It is just amazing how pretty is our Martian environment!

Around 18h00, we worked on our reports as well as on our own experiments.

Tonight, we plan to eat some potatoes with the rest of the vegetables and to play some board games.

This is it for today.

Martianly yours,
Tharsis commander

Sol Summary – March 30th

Crew 227 Sol Summary Report 30-03-22
Sol: 3

Summary Title: A successful EVA under inside instruction.
Author’s name: Jean Jacobs (XO)

Mission Status: The crew is getting more and more comfortable in its new environment.

Sol activity summary:

09:00 Breakfast and 3D printing of the new stove knob
10:00 Team “anatomy” and “mechanics” facing in Julie’s experiment
12:30 Lunch
13:15 Departure on EVA of 5 crew members
14:00 Cheyenne checked the environmental parameters in her experiment, Cyril calibrated his radar
16:00 Return from EVA crew
18:00 Report redaction and correction
19:00 Capcom
20:00 Dinner

Anomalies in work: None

Weather: Sunny day

Crew Physical Status: The crew is in shape

EVA: EVA to North Ridge

Reports to be filed:
– Crew commander
– Journalist
– GreenHab officer
– Engineer
– EVA report
– HSO
– Astronomer
– EVA request
– Some pictures
– Picture of the day
– Mission plan

Support Requested: None at this time

Sincerely,
Jean Jacobs

EVA Report – March 30th

Crew 227 EVA report 30/03/2022
EVA #4
Author: Jean Jacobs

Purpose of EVA: Placing Audrey’s cuvettes outside, placing Cyril’s antenna on Perseverance, and exploring North Ridge

Start time: 13:15 pm
End time: 16:00 pm

Narrative: The five crew members placed the Opportunity and Perseverance in front of the science dome window and transferred Cyril’s antenna from one to the other following his instructions. In the meanwhile, Jean and Audrey placed 10 cuvettes under Martian conditions. After that, the group went up to North Ridge and proceeded to an exploration of the surroundings finding what could be an interesting flat region for Cyril’s radar experiment.

Destination: North Ridge
Coordinates: 518500 E 4251500 N
Participants: Ignacio, Julie, Audrey, Julien, and Jean
Roads and routes per MDRS map: Cow Dung Road 0110
Mode of travel: Rovers Opportunity, Curiosity, and Spirit

Sincerely,
Jean Jacobs

Supplemental Operations Report – March 30th

Supplemental Operations Report 30 Mar 2022

Name of person filing report: Shannon Rupert

Reason for Report: Routine

Non-nominal systems: Nothing to report?…!!!

Action taken for non-nominal systems: n/a

Generator: Still being manually run at night.

ScienceDome Dual Split: Off

Solar—Charge controllers off at night.

Solar— Nominal

Notes on power system: Nothing to report

Propane Reading, station tank – 60%

Propane Reading, director tank— 62% (This is higher that my last reading)

Propane Reading, intern tank— 72%

Propane Reading, generator— 49%

Ethanol Free Gasoline – 0.5 gallons.

Water (Outpost tank) – 50 gallons

Sojourner rover used: no

Hours: 167.5

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Notes on rovers: There is no available distilled water in town.
ATV’s Used: (Honda, 350.1, 350.2, 300): none

Reason for use: n/a

Oil Added? No.

ATV Fuel Used: 0 gallons

# Hours the ATVs were Used: 0 hours

Notes on ATVs: Honda’s flat tire has not yet been taken for repair.

HabCar used and why, where? Yes, to town for garbage and water runs

CrewCar used and why, where? Driven to MDRS by the crew. Needs an oil change and tires rotated. I will go to Grand Junction next week to do these things.

Luna used and why, where? Yes, to town

Campus wide inspection, if action taken, what and why? Nothing to report.

General notes and comments: Nothing to report

Summary of general operations: Nothing to report

Summary of internet: Nothing to report

Summary of suits and radios: Nothing to report

Summary of Hab operations: The crew replaced one of the broken knobs on the stove with one that they 3D printed. It looks awesome. There are currently two broken knobs, so hopefully, they will be able to print another one during their time here.

Summary of Outpost operations: Nothing to report

Summary of GreenHab operations: The crew is using the GreenHab for space, but it is not being climate controlled.

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: I am concerned that the broken stair on the lower deck staircase may break completely by the end of the season. It will require welding repairs before next season.

Operations Report – March 30th

Crew 227 Operations Report 30-03-2022

SOL:3

Name of person filing report: Julien Meert

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report

ROVERS

Spirit rover used: Use for EVA
Hours: 164
Beginning charge: 100%
Ending charge: 92
Currently charging: Yes

Opportunity rover used: Yes
Hours: 91
Beginning charge:100%
Ending charge: 94
Currently charging: Yes

Curiosity rover used: Yes
Hours: 179
Beginning charge: 100%
Ending charge: 92
Currently charging: Yes

Perseverance rover used: No
Hours: I will check tomorrow
Beginning charge: 100%
Ending charge: Plugged in
Currently charging: Yes

General notes and comments: As you can see in the photo we repair the stove knob

Summary of Hab operations : Nothing to report

WATER USE:

Water (static tank): 388 gallons
Water (loft tank): 22 gallons
Water Meter: 1571300 units

Static to Loft Pump used: 27 gallons
Static tank pipe heater (on or off): Off
Static tank heater Off

Toilet tank emptied: No

Summary of internet: Stable

Summary of suits and radios: The ventilator of the one-piece switch seems not to work even charge.

Summary of GreenHab operations: Out of Service

WATER USE: 0 gallons
Heater: Off
Supplemental light: Off
Harvest: 0

Summary of ScienceDome operations: We put some plants on the front of the window
Dual split: Off

Summary of RAM operations: We try the Dremel yesterday but we were not able due to missing pieces. There is also a battery of the drill machine that is empty and the charger is also missing.

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing

Questions, concerns and requests to Mission Support: Not especially.

Sol Summary – March 25th

Crew 264 Sol Summary Report 3-25-22

Sol: 6

Summary Title: Discussion of medical futility

Author’s name: Arian Anderson

Mission Status: Tragedy struck

Sol Activity Summary: EVA took a turn when a crew member suffered from a simulated cardiac arrest. There were multiple simulated injured patients but the patient in the simulated arrest did not "survive". Everything possible was done and the crew is devastated.

Look Ahead Plan: Tonight is movie night and planning for evac in the am

Anomalies in work: None

Weather: Started getting hot

Crew Physical Status: Improving on idea of going home

EVA: See above

Reports to be file: Operations Report

Support Requested: None at this time.

Operations Report – March 25th

Crew 264 Operations Report 25-03-2022

SOL: 6

Name of person filing report: Arian Anderson

Non-nominal systems: None

Notes on non-nominal systems: None

ROVERS

Spirit rover used:

Hours: 162

Beginning charge: 100%

Ending charge: 75

Currently charging: No

Opportunity rover used:

Hours: 91

Beginning charge: 100

Ending charge: 76

Currently charging: No

Curiosity rover used:

Hours: 177

Beginning charge: 100

Ending charge: 73

Currently charging: No

Perseverance rover used:

Hours: 233

Beginning charge: 100

Ending charge: 70

Currently charging: No

General notes and comments: None

Summary of Hab operations:

WATER USE: 40 gallons (25 people)

Water (static tank): ~425 gallons

Water (loft tank): 90 gallons

Water Meter: 157230.4 units

Static to Loft Pump used – Yes

Static tank pipe heater (on or off): Off

Static tank heater (On or off): Off

Toilet tank emptied: Yes

Summary of internet: Connected to the internet

Summary of suits and radios: All working well

Summary of GreenHab operations: N/A not in use

WATER USE: gallons

Heater: On or Off

Supplemental light: On or Off

Harvest:

Summary of ScienceDome operations: None

Dual split: Heat or AC, On or Off

Summary of RAM operations: N/A

Summary of any observatory issues: N/A

Summary of health and safety issues: N/A.

Questions, concerns and requests to Mission Support: None